OpenStudy (anonymous):

Evaluate without using a calculator. Give answer in radians and degrees. Show work please or explain it!

OpenStudy (anonymous):

\[\cos^{-1}(-\sqrt{2}/2) \]

OpenStudy (dumbcow):

from the unit circle i know that cos(x) = -sqrt(2)/2 when x = 3pi/4 and 5pi/4

OpenStudy (dumbcow):

convert to degrees.....pi = 180

OpenStudy (dumbcow):

do you need more explanation?

OpenStudy (anonymous):

thank you so its 180 degrees & 3 pi/4 ?

OpenStudy (dumbcow):

?? no pi radians is equivalent to 180 degrees there are 2 possible answers (2 diff angles satisfy cos(x) = -sqrt2/2) i gave you answers in radians, you have to convert to degrees by replacing pi with 180

OpenStudy (anonymous):

oh... i'm sorry so the answer is 180 degrees & pi ? I don't understand how to get to 180 degrees. :/

OpenStudy (dumbcow):

hmm ok apparently i am only confusing you more just look at the unit circle and find -sqrt2/2, the answers will be the angle measures

OpenStudy (anonymous):

5 pi/4 ?

OpenStudy (dumbcow):

yep thats 1 answer

OpenStudy (anonymous):

i hope i'm not annoying or anything lol you don't have to help me but is the answer for \[\sin^{-1} (-1)\] = 270 degrees, (3 pi/2) ?